### v1

# Test-Time Adaptation of Vision-Language Models for Open-Vocabulary Semantic Segmentation

Abstract
Recently, test-time adaptation has attracted wide interest in the context of vision-language models for image classification. However, to the best of our knowledge, the problem is completely overlooked in dense prediction tasks such as Open-Vocabulary Semantic Segmentation (OVSS). In response, we propose a novel TTA method tailored to adapting VLMs for segmentation during test time. Unlike TTA methods for image classification, our Multi-Level and Multi-Prompt (MLMP) entropy minimization integrates features from intermediate vision-encoder layers and is performed with different text-prompt templates at both the global CLS token and local pixel-wise levels. Our approach could be used as plug-and-play for any segmentation network, does not require additional training data or labels, and remains effective even with a single test sample. Furthermore, we introduce a comprehensive OVSS TTA benchmark suite, which integrates a rigorous evaluation protocol, seven segmentation datasets, and 15 common corruptions, with a total of 82 distinct test scenarios, establishing a standardized and comprehensive testbed for future TTA research in open-vocabulary segmentation. Our experiments on this suite demonstrate that our segmentation-tailored method consistently delivers significant gains over direct adoption of TTA classification baselines. Our code can be found at our repository.
1 Introduction
Contrastive Vision-Language Models (VLMs) such as CLIP clip have demonstrated remarkable generalization capabilities by aligning vision and language modalities through large-scale pre-training. This versatility has positioned VLMs as powerful foundation models for numerous downstream tasks lin2022frozen ; guzhov2022audioclip ; liu2023clip . A promising direction for leveraging VLMs beyond classification is Open-Vocabulary Semantic Segmentation (OVSS), where models aim to segment objects beyond a pre-defined set of categories, via VLMs’ zero-shot recognition capabilities. Unlike traditional segmentation methods that require pixel-wise supervision, OVSS enables generalization to unseen object categories through language-driven representations.
Although existing OVSS methods have made significant progress, they remain vulnerable to domain shifts at test time, such as environmental changes or image corruptions, which may dramatically degrade segmentation quality. In the absence of a mechanism enabling to adapt to unseen test-time distributions, these models might lose their generalization capabilities, which limits their reliability in real-world applications. Consequently, there is an unresolved gap for the Test-Time Adaptation (TTA) of OVSS models, which would enable models to dynamically adjust both to the task shift of VLM-based segmentation and to the domain shifts encountered during inference.
To close this gap, we present a novel Multi-Level Multi-Prompt (MLMP) test time adaptation strategy, the first fully test-time adaptation framework that could be plugged into any OVSS model, to the best of our knowledge. MLMP is lightweight and plug-and-play, boosting performance on the fly without access to labels. Its power comes from two key ideas: (i) adaptively integrating intermediate vision-encoder layers to harvest complementary, shift-resilient features, and (ii) a multi-prompt optimization that exploits VLMs’ template sensitivity to provide a robust adaptation signal across diverse text-template conditions.
The core requirement for test-time adaptation is a reliable signal that faithfully reflects the current input distribution—even under severe domain shifts or corruptions. To meet this need, MLMP begins by adaptively integrating intermediate layers of the vision encoder: earlier layers preserve fine-grained edges and textures, while deeper blocks encode semantic context, and each layer reacts differently when the data distribution changes. By aggregating these multi-level features into the adaptation process and weighting them by their confidence, MLMP harvests the most trustworthy signals for each input sample.
Beyond multi-level fusion, MLMP leverages VLMs’ prompt sensitivity to model uncertainty. Prior work watt ; tpt shows that changing a prompt template, e.g., from ‘’a photo of a {class}” to ‘’an origami of a {class}”, could drastically change the classification performance. In segmentation, we show that this effect is even more extreme: per-pixel predictions under different prompts diverge far more than the single CLS token used for classification (Fig. 1a: CLS token vs. last-layer spatial tokens). This sensitivity effect is even more pronounced in intermediate feature maps. The intermediate layers that we fuse for reliability exhibit even stronger template-specific shifts (Fig. 1a: intermediate-layer tokens). Instead of viewing this inconsistency as a weakness, MLMP models it directly by incorporating multi-prompt, multi-level predictions into its adaptation objective function. This multi-prompt approach not only smooths out template-specific noise, thereby reducing gradient variance and preventing degenerate collapse, but also ensures that the model yields segmentations that are consistent across diverse linguistic formulations. In this way, MLMP transforms prompt sensitivity into a powerful adaptation signal that complements its multi-level feature integration. As illustrated in Fig. 1b, our MLMP method consistently outperforms the non-adapted baseline, achieving about 8–9 absolute mIoU improvements on domain-shifted inputs, while also boosting performance on original (non-corrupted) images. This demonstrates the benefit of our joint multi-level, multi-prompt adaptation.
Figure 1: Motivation. (a) Left: Mean $\pm$ std entropy across seven text templates for the CLS token and the spatial tokens of the final and intermediate vision layers. Even the final-layer spatial tokens exhibit higher entropy and variability than CLS, and this sensitivity grows further in intermediate layers (numbers show % std increase relative to CLS). These patterns highlight pronounced prompt-induced uncertainty at multiple depths and motivate both multi-level and multi-prompt adaptation. (b) Right: mIoU of the baseline vs MLMP on clean and corrupted data, showing consistent absolute improvements and underscoring the effectiveness of our joint adaptation strategies.
We outline our key contributions as follows:
  • Plug-and-Play TTA Framework for OVSS: We introduce MLMP, which is, to the best of our knowledge, the first fully test-time adaptation method that could be easily applied to any OVSS backbone.
  • Adaptive Multi-Level Fusion: MLMP integrates features from intermediate vision-encoder layers to capture complementary, shift-resilient cues. To further enhance robustness, we propose an uncertainty-aware strategy that re-weights features from individual layers based on their prediction entropy.
  • Multi-Prompt Local-Global Test Time Optimization: MLMP turns prompt sensitivity into signal by directly minimizing entropy across different text prompt templates at both the global CLS token and local pixel-wise levels. This optimization naturally complements our mutli-layer feature fusion by enforcing consistency across linguistic perspectives and feature depths.
  • Comprehensive OVSS TTA Benchmark Suite: We curate a rigorous evaluation protocol spanning seven mainstream segmentation datasets and 15 common corruptions, with a total of 82 distinct test scenarios, establishing a standardized and comprehensive testbed for future TTA research in open-vocabulary segmentation. Our experiments on this suite demonstrate that MLMP consistently delivers significant gains over baselines across all scenarios.
2 Related Work
Test-time adaptation (TTA) for open-vocabulary semantic segmentation (OVSS) remains unexplored—existing TTA methods focus on classification or single-modal segmentation, while OVSS approaches use VLMs without any online adaptation. We bridge this gap with our proposed method MLMP, a plug-and-play TTA framework that could be applied to any OVSS method.
Test-Time Adaptation.  TTA addresses domain shifts by adapting pre-trained models to unlabeled target data without source samples. Methods like PTBN PTBN and TENT tent update batch statistics and affine parameters via entropy minimization but rely on large batches or augmentations. MEMO memo simplifies this with single-sample augmentations, LAME lame clusters features via Laplacian smoothing, and SAR SAR stabilizes adaptation using batch-agnostic normalization and sharpness-aware entropy minimization.
Test-Time Adaptation on Segmentation.  TTA enhances segmentation robustness against domain shifts without source data. Methods include self-supervised adaptation via entropy minimization or contrastive learning tent ; liu2021source , single-image adaptation optimizing per-image predictions he2021autoencoder , and continual TTA that leverages clustering to prevent forgetting mummadi2021test . Multi-modal adaptation uses cross-modal self-supervision valvano2023learning , while active TTA integrates minimal human feedback for guided refinement wang2023active . These approaches assume a fixed, vision-only label space and rely on spatial or surrogate tasks, making them ill-suited for zero-shot, text-driven OVSS. Consequently, none have been applied to VLMs.
Open-Vocabulary Semantic Segmentation.  OVSS enables segmentation of unseen categories using vision-language models like CLIP. Approaches fall into fully-supervised, weakly-supervised, and training-free categories. Fully-supervised methods use pixel-wise annotations barsellotti2023enhancing ; ghiasi2022scaling ; liang2023open , while weakly-supervised ones leverage image-text pairs cai2023mixreorg ; cha2023learning ; chen2023exploring ; xu2022groupvit . Training-free OVSS avoids adaptation data but may rely on auxiliary pre-trained models barsellotti2024fossil ; corradini2024freeseg ; karazija2023diffusion . Training-free OVSS approaches aim to enhance segmentation without additional training data. Some methods, such as SCLIP wang2025sclip , adjust self-attention mechanisms to improve feature localization, while others, like MaskCLIP zhou2022extract , refine feature extraction from CLIP’s visual backbone. GEM bousselham2024grounding introduces additional optimization techniques to extract better dense features without fine-tuning. Among these, NACLIP naclip enhances CLIP’s dense prediction capabilities by introducing neighborhood attention, which ensures that image patches focus on nearby regions, and by refining similarity measures to improve spatial consistency.
To the best of our knowledge, this is the first work to address TTA for OVSS models, filling a previously unexplored intersection between these fields.
Figure 2: Overview of our MLMP method. In the Adaptation Phase, the model is adapted by leveraging multiple prompt templates alongside various intermediate feature layers, as well as the global feature. During the Evaluation Phase, the model computes weights based on the entropy of the intermediate features to perform a weighted averaging. These averaged features, combined with the different templates, are then used to generate the final segmentation map.
3 Methodology
We first revisit the contrastive vision–language model (VLM) for open-vocabulary semantic segmentation (OVSS), and then present our Multi-Level Multi-Prompt (MLMP) adaptation strategy.
3.1 OVSS with VLMs
Given an input image $\mathbf{X}\in{}^{H\times W\times 3}$ and a set of concepts $C_{k}\in\mathcal{C}$ expressed in natural language, OVSS seeks a semantic mask $\mathbf{y}\!\in\!\{1,\dots,K\}^{H\times W}$ that assigns one concept to every pixel.
Following recent approaches for OVSS naclip ; wang2025sclip , we employ a transformer-based VLM to extract visual and text features from the image and concepts in natural language. Specifically, we feed the image $\mathbf{X}$ into the ViT-based vision encoder to extract a visual token matrix $\mathbf{F}=\big{[}\mathbf{f}_{[\texttt{cls}]},\mathbf{f}_{1},\dots,\mathbf{f}_{N}\big{]}$ with each $\mathbf{f}_{i}\in{}^{D}$, where $N\!=\!\lfloor H/s\rfloor\!\times\!\lfloor W/s\rfloor$ is the number of patches of size $s\!\times\!s$ in the image and [cls] is the CLS token for classification. We define $\mathbf{Q}=\big{[}\mathbf{q}_{[\texttt{cls}]},\mathbf{q}_{1},\dots,\mathbf{q}_{N}\big{]}$, with each $\mathbf{q}_{i}\in{}^{D^{\prime}}$, the output features before the projection layer: $\mathbf{F}=\text{proj}(\mathbf{Q})$. At the same time, the text encoder is employed to extract text features $\mathbf{t}_{k}\in{}^{D}$ for each concept $C_{k}\in\mathcal{C}$. This is achieved by combining $C_{k}$ with a text prompt template, for instance “A photo of a [$C_{k}$]” or “An image of a [$C_{k}$]” where $C_{k}$ is an arbitrary text description like “white horse”.
The standard approach for classifying images with a contrastive VLM such as CLIP clip computes the cosine between the CLS token features and text embeddings of classes, and assigns the image to the class with highest similarity:
[…]
For extending this approach to segmentation, we instead compute the similarity between patch embeddings $\mathbf{f}_{i}$ and text embeddings $\mathbf{t}_{k}$ and assign a class/concept to each patch.
3.2 MLMP: Proposed Method
Figure 2 illustrates our full test-time adaptation pipeline, MLMP. MLMP integrates three complementary ideas: uncertainty-aware multi-level fusion, image-level entropy minimization and multi-prompt adaptation.
We begin by modifying the entropy minimization objective of TENT tent from image classification to work with spatial tokens. More specifically, for a batch of $B$ images, each containing $N$ tokens, the probability that token $i$ belongs to concept $k$ is
[…]
where $\tau$ is a softmax temperature scaling parameter, $\mathbf{T}=\big{[}\mathbf{t}_{1},\ldots,\mathbf{t}_{|\mathcal{C}|}\big{]}$, and $\mathrm{norm}(\cdot)$ denotes a function normalizing the columns of its input matrix to unit length. Also, let $\mathbf{P}$ denote a matrix containing the probabilities in (2), which could be expressed more compactly as follows:
$\mathbf{P}\,=\,\text{softmax}\big{(}\mathrm{norm}(\mathbf{F})\!\cdot\!{\mathrm{norm}(\mathbf{T})^{\top}}/\tau\big{)}.$ (3)
The batch-wise entropy, which is minimized for adaptation, is then defined as follows:
$\mathcal{H}(\mathbf{P})\,=\,-\frac{1}{B\!\cdot\!N}\sum_{i=1}^{B\cdot N}\sum_{k=1}^{|\mathcal{C}|}p_{ik}\log p_{ik}.$ (4)
Following watt ; clipartt , we keep the entire text encoder frozen and update only the LayerNorm parameters of the vision encoder during adaptation. Freezing the text encoder greatly reduces computational overhead, since text embeddings can be precomputed and reused across all test samples.
Uncertainty-Aware Multi-Level Fusion.  In VLM-based classification, the CLS token in the last layer of the visual encoder is typically used to compute the class label probabilities. This approach relies on the idea that the relevant information for classification lies at the end of the ViT and that intermediate layers serve to transform features. In segmentation, however, features from intermediate layers are often used to capture complementary information at different scales unet . This hypothesis is validated in Fig. 1b, and more specifically in Table1), showing that a higher segmentation mIoU is obtained when combining the features from different layers near the end of the ViT.
Inspired by this result, and leveraging the useful property of ViTs that the output of each layer has the same shape, we extend the entropy-based loss described above to use features from multiple layers. Denoting as $\mathbf{q}^{\ell}_{i}$ the visual features of patch $i$ obtained at layer $\ell$, we seek to aggregate the multi-level features into a single vector $\overline{\mathbf{q}}_{i}$ for segmentation prediction. A simple approach for doing this is to compute $\overline{\mathbf{q}}_{i}$ by averaging $\mathbf{q}^{\ell}_{i}$ across all layers $\ell$. However, this approach ignores the relative contribution and confidence of each layer in the final segmentation. To address this limitation, we estimate a confidence weight $\alpha^{\ell}$ for each layer $\ell$ based on its prediction entropy. First, we get visual features $\mathbf{F}^{\ell}_{i}=\text{proj}(\mathbf{Q}^{\ell}_{i})$ using the same projection head as for the final segmentation. Following the same approach as before, we then compute the batch-wise entropy of layer $\ell$ as
$h^{\ell}\,=\,\mathcal{H}(\mathbf{P}^{\ell}),\ \ \text{with }\mathbf{P}^{\ell}=\text{softmax}\big{(}\mathrm{norm}(\mathbf{F}^{\ell})\cdot{\mathrm{norm}(\mathbf{T})^{\top}}/\tau\big{)}.$ (5)
Finally, the confidence weight of the layer is obtained using a softmax as follows:
$\alpha^{\ell}\,=\,\frac{\exp(-\beta\!\cdot\!h^{\ell})}{\sum_{\ell^{\prime}=1}^{L}\exp(-\beta\!\cdot\!h^{\ell^{\prime}})}.$ (6)
Here, $\beta$ is a parameter controlling the “sharpness” of the weight distribution. During adaptation, we set $\beta=0$ to promote a uniform contribution from all layers in the prediction. During inference, we sharpen the distribution with a value of $\beta=1$, emphasizing the more confident layers in the final prediction.
With these confidence weights, we can now obtain our uncertainty-aware multi-level (UAML) features as
$\overline{\mathbf{F}}\,=\,\text{proj}(\overline{\mathbf{Q}}),\ \ \text{with }\overline{\mathbf{Q}}=\sum_{\ell=1}^{L}\alpha^{\ell}\mathbf{Q}^{\ell},$ (7)
giving the following entropy-based loss to minimize:
$\mathcal{L}_{\mathrm{UAML}}(\mathbf{T})\,=\,\mathcal{H}(\overline{\mathbf{P}}),\ \ \text{with }\overline{\mathbf{P}}=\text{softmax}\big{(}\mathrm{norm}(\overline{\mathbf{F}})\cdot{\mathrm{norm}(\mathbf{T})^{\top}}/\tau\big{)}.$ (8)
Image-Level Entropy Minimization.  Since the CLS token is not directly linked to individual patch predictions but rather captures a more global representation of the input, we also include an image-level entropy (ILE) minimization term specifically for this token. As illustrated in Fig. 1, the CLS token demonstrates increased robustness and reliability. This term, which encourages the model to produce more confident global predictions is expressed as:
$\mathcal{L}_{\mathrm{ILE}}(\mathbf{T})\,=\,-\frac{1}{B}\sum_{b=1}^{B}\sum_{k=1}^{|\mathcal{C}|}p_{b,k}^{[\texttt{cls}]}\log p_{b,k}^{[\texttt{cls}]}.$ (9)
Here, $p_{b,k}^{[\texttt{cls}]}$ denotes the predicted probability for concept $C_{k}$ obtained using the CLS token in the last layer for the $b$-th sample.
Multi-Prompt Adaptation.  Prior work on TTA for classification watt has shown the usefulness of leveraging multiple prompt templates in VLM to encode class labels, based on the idea that the templates capture complementary information about these classes. As shown in Fig. 1a, the sensitivity to the choice of prompt templates is even more pronounced in segmentation tasks, where fine-grained spatial predictions are required. Using multiple templates acts as cross-modal regularization, encouraging more stable and generalized learning signals. While different from image augmentation, it can be seen as a strong, safe, and lightweight text-space augmentation. Rather than averaging the weights adapted from different prompt templates as in watt —a computationally expensive approach for dense prediction tasks such as segmentation—our method minimizes our proposed UAML and ILE losses across these templates. Let $\mathbf{T}_{t}$ be the text features obtained using the $t$-th template. Our final adaptation loss is defined as:
$\mathcal{L}_{\mathrm{final}}(\theta)\,=\,\frac{1}{T}\sum_{t=1}^{T}(\mathcal{L}_{\mathrm{UAML}}(\mathbf{T}_{t})+\mathcal{L}_{\mathrm{ILE}}(\mathbf{T}_{t})).$ (10)
Theoretical Justification.  Each template $t$ contributes its own adaptation loss, as we optimize their average in Eq. (10). By optimizing the adaptation loss of each prompt directly, we force the model to correct for the unique wording and visual cue of each template, rather than ’averaging’ these differences in the text embedding space. This loss-level integration treats each template as an independent critic, translating diverse linguistic perspectives into separate gradient signals. Averaging those signals produces an unbiased descent direction whose variance decays as $1/T$, enabling each adaptation step to represent the full prompt ensemble while being stable under noisy shifts.
Proposition 1 (Unbiasedness and Variance Bound).
Assume that each per-template gradient $g_{t}(\theta)=\nabla_{\theta}\bigl{[}\mathcal{L}_{\mathrm{UAML}}(\mathbf{T}_{t})+\mathcal{L}_{\mathrm{ILE}}(\mathbf{T}_{t})\bigr{]}$ has variance bounded by $\sigma^{2}$, then the ensemble gradient, defined by $\nabla_{\theta}\mathcal{L}_{\mathrm{final}}=\frac{1}{T}\sum_{t=1}^{T}g_{t}(\theta)$, is unbiased and satisfies the following variance bound:
$\mathbb{E}\Bigl{[}\nabla_{\theta}\mathcal{L}_{\mathrm{final}}\Bigr{]}=\mathbb{E}\bigl{[}g_{t}(\theta)\bigr{]};\quad\operatorname{Var}\!\Bigl{(}\nabla_{\theta}\mathcal{L}_{\mathrm{final}}\Bigr{)}=\frac{1}{T^{2}}\sum_{t=1}^{T}\operatorname{Var}\bigl{(}g_{t}(\theta)\bigr{)}\leq\frac{\sigma^{2}}{T}$ (11)
Proof.
A proof of Prop. 1 is provided in Appendix A. ∎
The $1/T$ reduction in gradient variance, as stated in Prop. 1, explains the improved stability we observe. Table 5 confirms that this loss-level ensemble outperforms alternative fusion strategies.
4 Experimental Settings
Experimental Setup.  Following prior work on TTA in classification clipartt ; watt , we restrict updates to the normalization layers within the vision encoder. The adaptation process is carried out over $10$ iterations using the Adam optimizer with a constant learning rate of $10^{-3}$ across all datasets. We use a batch size of $2$ images during adaptation across all datasets. For each new batch, the model undergoes a reset, restoring it to its initial weights before adaptation is applied.
Datasets.  In traditional TTA for segmentation, two datasets are commonly employed to simulate domain shifts—one for model training and another for adaptation during inference (e.g., GTAV and Cityscapes)—as both must share the same semantic label space. In our study, as this is the first exploration of TTA for VLMs in segmentation tasks and given that VLMs are pre-trained, we draw inspiration from ImageNet-C cifar10c&10.1 to introduce 15 synthetic corruptions on segmentation datasets. Our experiments are conducted on Pascal VOC 20 (v20), Pascal VOC 21 (v21) PascalVoc , Pascal Context 59 (P59), Pascal Context 60 (P60) PascalContext , and Cityscapes Cityscapes , incorporating both original version (clean) and the synthetic 15 corruptions (denoted with a “-C” suffix). For COCO-Stuff CocoStuff and COCO-Object CocoObject , we use only the original versions. This setup results in 82 distinct test scenarios covering both synthetic and realistic domain shifts, with variations in resolution, scene diversity, object sizes, and class granularity—enabling a comprehensive evaluation of our method’s robustness.
Benchmarking.  While MLMP is compatible with any OVSS framework, we incorporate NACLIP naclip with ViT-L/14 as our baseline OVSS model, which leverages neighborhood attention to enhance spatial consistency in a training-free manner. The compared methods include TENT tent , which serves as a baseline and minimizes entropy during adaptation; CLIPArTT clipartt , which employs pseudo-labels generated via conformal learning; WATT watt , which averages learnable parameters across multiple parallel branches; and TPT tpt , which performs prompt tuning to adapt VLMs at test time. For a fair comparison, we modified all methods for the segmentation setting by processing all spatial tokens extracted from the VLM, rather than relying solely on the CLS token.
Table 1: mIoU performance when using different layer ranges in the proposed multi-level adaptation.
ViT-L/14 Layer Range $L_{24}$ (last) $L_{23\text{--}24}$ (last two) $L_{22\text{--}24}$ (last three) $L_{19\text{--}24}$ (last 25%) $L_{13\text{--}24}$ (last 50%) $L_{7\text{--}24}$ (last 75%) $L_{1\text{--}24}$ (all layers)
V20 (Original) 77.00 $\pm$0.04 77.65 $\pm$0.02 77.66 $\pm$0.09 80.61 $\pm$0.05 80.50 $\pm$0.03 81.67 $\pm$0.04 78.79 $\pm$0.02
Gaussian Noise 63.02 $\pm$0.06 64.41 $\pm$0.05 65.39 $\pm$0.13 66.88 $\pm$0.18 66.88 $\pm$0.02 67.82 $\pm$0.01 63.06 $\pm$0.09
Defocus Blur 72.06 $\pm$0.12 72.93 $\pm$0.19 72.84 $\pm$0.02 76.10 $\pm$0.16 76.37 $\pm$0.05 78.78 $\pm$0.02 77.56 $\pm$0.09
Snow 71.04 $\pm$0.05 72.09 $\pm$0.04 72.56 $\pm$0.02 74.47 $\pm$0.12 74.41 $\pm$0.01 76.39 $\pm$0.02 73.72 $\pm$0.07
JPEG Compression 71.84 $\pm$0.15 73.88 $\pm$0.11 74.40 $\pm$0.07 76.96 $\pm$0.02 77.67 $\pm$0.03 78.73 $\pm$0.08 75.87 $\pm$0.19
V20-C Average 69.33 70.33 70.78 72.89 73.45 74.90 72.02
5 Results
5.1 Ablation studies
Effect of Intermediate Layers.  To analyze the impact of layer selection in our uncertainty-aware multi-level adaptation strategy, we use different ranges of intermediate layers. As shown in Table 1, performance varies notably with the fusion range. While using only the final layers yields moderate improvements, incorporating the last 75% of the layers consistently achieves the best performance across both clean and corrupted inputs. This highlights that multi-level fusion is a key driver of adaptation performance: earlier layers, although less semantically abstract, contribute valuable low-level features—such as texture and edge cues—that enhance robustness to distribution shifts. In this ablation, we isolate the effect of multi-level fusion by applying only the first term in Eq. 10, using a single prompt template and omitting the $L_{ILE}$ term.
Effect of Uncertainty-Aware Layer Fusion.  We investigate strategies for aggregating multi-level features by comparing uniform averaging ($\beta=0$) and uncertainty-aware fusion ($\beta=1$) during evaluation, using the same 75% layer range identified in the previous ablation. As shown in Table 2, incorporating entropy-based weighting improves performance by 4.29% on V20 and 3.31% on V20-C. This highlights the importance of leveraging layer-wise confidence when aggregating features.
Figure 3: Mean and standard deviation of layer-wise confidence weights of MLMP across datasets. The fusion mechanism adaptively emphasizes more reliable layers based on input conditions.
Visualization of Layer-Wise Confidence Weights.  We visualize the mean and standard deviation of the learned layer weights to better understand the behavior of our uncertainty-aware fusion strategy. As shown in Fig. 3, deeper layers tend to receive higher confidence, though earlier layers also contribute, especially under corrupted conditions. This variation is most pronounced in the Cityscapes dataset, where the distribution fluctuates more across layers and corruption types. In contrast, the COCO-Stuff dataset shows a flatter distribution, where the final layer is not consistently the most influential. These results underscore the core strength of our fusion mechanism: its ability to adaptively reweight layers based on input conditions, assigning greater importance to those that remain more reliable under distribution shifts and corruption. Please refer to Appendix H for additional results on other datasets.
Effect of Global Image-Level Adaptation Term.  The image-level entropy term, $L_{ILE}$ complements our patch-level adaptation by encouraging consistent global predictions through the CLS token. While the multi-level loss targets fine-grained spatial predictions, the ILE term introduces global context that helps stabilize adaptation. As shown in Table 2, when added in isolation, $L_{ILE}$ improves performance by 1.74% and 2.66% on V20 and V20-C, respectively, demonstrating the benefit of incorporating global context under distribution shift.
Table 2: mIoU comparison of MLMP components, showing individual and combined contributions.
Multi-Level Fusion
Multi-Prompt Loss
Image-Level Entropy
Uncertainty-Aware Weighting
V20 (Original) 77.00 77.38 81.67 79.70 78.74 78.97 83.00 77.69 82.70 81.15 79.13 83.76
Gaussian Noise 63.02 65.42 67.82 66.75 65.66 65.96 69.13 66.17 69.00 69.62 67.35 71.13
Defocus Blur 72.06 76.65 78.78 74.31 75.00 76.46 78.78 77.29 79.78 77.14 77.79 80.36
Snow 71.04 72.64 76.39 74.66 74.16 73.25 77.31 74.05 78.50 77.20 74.94 79.53
JPEG Compression 71.84 74.38 78.73 75.56 74.77 76.77 80.81 74.61 79.79 77.98 77.94 82.06
V20-C 69.33 71.59 74.90 72.58 71.99 72.66 75.97 72.41 76.18 75.08 73.89 77.58
Figure 4: mIoU performance of our method for different numbers of templates.
Figure 5: mIoU performance for prompt-integration strategies (Text, Params, Loss) on clean and corrupted data.
Dataset: V20 Text Params Loss
Original 78.91 $\pm$0.07 74.46 $\pm$0.21 79.70 $\pm$0.06
Gaussian Noise 66.27 $\pm$0.00 62.83 $\pm$0.04 66.75 $\pm$0.01
Defocus Blur 74.05 $\pm$0.10 70.28 $\pm$0.16 74.31 $\pm$0.09
Snow 73.78 $\pm$0.02 70.10 $\pm$0.30 74.66 $\pm$0.01
JPEG Compression 74.98 $\pm$0.05 70.55 $\pm$0.11 75.56 $\pm$0.02
V20-C Average 71.92 68.44 72.58
Effect of Number of Prompt Templates.  To isolate this effect, we evaluate performance using different numbers of prompt templates while disabling both the multi-level fusion and the image-level entropy (ILE) term in Eq.10. As shown in Fig.5, increasing the number of templates improves performance up to 7, after which the gains begin to saturate or slightly decline. This trend holds across both clean and corrupted settings, indicating that a moderate number of diverse prompt templates is sufficient for MLMP. We therefore use 7 templates by default in our main experiments. Please refer to Appendix C for template details.
Where to Integrate Multi-Prompt Information.  Here, we empirically compare strategies for integrating multi-prompt information into the adaptation process. Specifically, we evaluate: (1) text-level averaging, where prompt embeddings are averaged before computing logits—a technique commonly used in zero-shot learning clip ; (2) a learnable parameter averaging baseline (Params) inspired by WATT watt ; and (3) our proposed method (Loss), which incorporates all prompt templates directly into the adaptation loss (Eq.10). To isolate the effect of prompt integration, this analysis excludes other components such as multi-level fusion and the image-level entropy (ILE) term. As shown in Table 5, our loss-level formulation consistently outperforms the alternatives across both clean and corrupted settings.
Full Component Analysis.  Table 2 presents an extensive ablation evaluating the contribution of each component in our MLMP strategy. Each proposed element yields consistent gains when added independently, but it is their combination that delivers the highest overall performance across both clean and corrupted settings, highlighting their strong complementary effects within a unified framework. Additionally, we provide further ablations in the Appendix, including alternative OVSS backbones, a different VLM, ViT architecture variants, computational complexity analysis, and MLMP segmentation map visualizations.
Table 3: mIoU comparison of MLMP and baselines across several datasets. CLIPArTT could not be run for a few cases owing to GPU memory shortages. Full per-dataset results are in Appendix J.
OVSS Backbone: NACLIP Adaptation Method
Dataset No Adapt. TENT TPT WATT CLIPArTT MLMP
V20 (Original) 75.91 77.00 $\pm$0.04 75.93 $\pm$0.01 57.73 $\pm$0.06 72.77 $\pm$0.14 83.76 $\pm$0.00
V20-C Gaussian Noise 62.89 63.02 $\pm$0.06 62.98 $\pm$0.01 36.44 $\pm$0.04 53.36 $\pm$0.25 71.13 $\pm$0.09
Shot noise 66.26 65.88 $\pm$0.06 66.33 $\pm$0.02 40.95 $\pm$0.05 58.15 $\pm$0.28 75.02 $\pm$0.03
Impulse Noise 63.16 64.17 $\pm$0.04 63.12 $\pm$0.01 34.90 $\pm$0.06 54.83 $\pm$0.03 71.34 $\pm$0.11
Defocus blur 72.59 72.06 $\pm$0.12 72.55 $\pm$0.02 52.43 $\pm$0.03 65.39 $\pm$0.45 80.36 $\pm$0.06
Glass blur 71.44 70.74 $\pm$0.07 71.40 $\pm$0.01 49.96 $\pm$0.05 64.62 $\pm$0.13 78.84 $\pm$0.05
Motion blur 73.10 73.50 $\pm$0.10 73.16 $\pm$0.02 53.35 $\pm$0.06 67.48 $\pm$0.17 81.41 $\pm$0.05
Zoom blur 59.03 61.36 $\pm$0.07 59.00 $\pm$0.01 41.39 $\pm$0.08 52.37 $\pm$0.12 69.41 $\pm$0.12
Snow 71.49 71.04 $\pm$0.05 71.44 $\pm$0.01 51.18 $\pm$0.06 66.97 $\pm$0.02 79.53 $\pm$0.05
Frost 65.38 67.01 $\pm$0.02 65.46 $\pm$0.01 45.75 $\pm$0.05 60.48 $\pm$0.08 73.20 $\pm$0.07
Fog 70.69 70.54 $\pm$0.07 70.70 $\pm$0.01 52.96 $\pm$0.04 67.85 $\pm$0.10 79.81 $\pm$0.06
Brightness 74.95 75.61 $\pm$0.02 74.95 $\pm$0.01 55.82 $\pm$0.05 71.52 $\pm$0.14 83.51 $\pm$0.01
Contrast 71.51 70.51 $\pm$0.04 71.49 $\pm$0.02 50.74 $\pm$0.06 66.01 $\pm$0.06 79.06 $\pm$0.16
Elastic transform 62.86 65.78 $\pm$0.05 62.95 $\pm$0.01 45.45 $\pm$0.04 60.41 $\pm$0.10 74.03 $\pm$0.01
Pixelate 77.28 76.95 $\pm$0.12 77.31 $\pm$0.01 59.76 $\pm$0.05 73.14 $\pm$0.17 84.97 $\pm$0.04
JPEG compression 72.59 71.84 $\pm$0.15 72.56 $\pm$0.01 53.44 $\pm$0.05 68.21 $\pm$0.07 82.06 $\pm$0.01
Mean 69.01 69.33 69.03 48.30 63.39 77.58
V21 (Original) 45.12 45.65 $\pm$0.02 45.17 $\pm$0.01 28.58 $\pm$0.05 39.50 $\pm$0.04 50.78 $\pm$0.02
V21-C 40.75 40.95 40.77 24.12 34.16 46.25
P59 (Original) 28.23 28.73 $\pm$0.02 28.26 $\pm$0.01 16.55 $\pm$0.04 24.60 $\pm$0.03 31.95 $\pm$0.02
P59-C 23.88 23.88 23.88 13.37 19.72 27.03
P60 (Original) 24.95 25.29 $\pm$0.01 24.98 $\pm$0.01 14.77 $\pm$0.03 21.88 $\pm$0.03 27.99 $\pm$0.03
P60-C 21.39 21.25 21.49 12.08 17.79 24.07
CityScapes (Original) 29.49 30.54 $\pm$0.04 29.57 $\pm$0.01 20.77 $\pm$0.06 33.35 $\pm$0.03
CityScapes-C 21.63 21.64 21.60 13.45 23.02
COCOObject (Original) 23.80 24.88 $\pm$0.01 23.84 $\pm$0.01 14.14 $\pm$0.06 21.34 $\pm$0.03 28.84 $\pm$0.01
COCOStuff (Original) 18.34 18.76 $\pm$0.01 18.35 $\pm$0.01 9.49 $\pm$0.02 15.48 $\pm$0.01 21.25 $\pm$0.01
5.2 Final Comparison with Alternative Adaptation Methods
Performance on Clean Data (No Distributional Shift).  We begin by evaluating MLMP on clean test data (original), where no distributional shift is present. This setting is crucial, as TTA methods must avoid degrading performance when adaptation is unnecessary. As shown in Table 3, MLMP achieves strong mIoU gains of +7.85, +5.66, +3.72, and +3.04 on V20, V21, P59, and P60, and +5.04/+2.91 on challenging datasets COCOObject and COCOStuff. In contrast, most alternative adaptation methods fail to improve performance in this setting. These gains highlight the robustness and generalization of our method, even when no explicit domain shift is present.
Performance Under Distributional Shift.  Under distributional shifts, the advantages of MLMP become even more apparent. As shown in Table 3, MLMP consistently outperforms both the zero-shot baseline and existing adaptation methods, achieving mIoU gains of +8.60, +5.50, +3.15, and +2.68 on V20-C, V21-C, P59-C, and P60-C, respectively. Beyond standard corruptions, we further evaluate on the Cityscapes dataset, which presents natural domain shifts such as environmental variation, weather conditions, and resolution differences. Despite its challenging nature and low zero-shot performance, MLMP improves mIoU by +3.86, demonstrating its real-world adaptability. To push this further, we apply 15 corruption types to create Cityscapes-C, where MLMP still yields a +1.39 gain. While TENT provides modest improvements, most other adaptation methods—including ClipArTT, TPT, and WATT—either fail to improve or degrade performance. These results highlight that naive strategies like pseudo-labeling, prompt tuning, or weight averaging are insufficient for open-vocabulary segmentation, and emphasize the need for segmentation-specific adaptation techniques such as MLMP. Detailed results can be found in Appendix H.
6 Conclusion
We presented MLMP, a plug-and-play test-time adaptation framework for open-vocabulary semantic segmentation that can be integrated with any OVSS method. By combining uncertainty-aware fusion of intermediate ViT features with a novel loss-level integration of multiple prompt templates, MLMP consistently enhances performance across both clean and shifted domains—including common corruptions and natural distributional shifts. Our comprehensive OVSS-TTA benchmark—covering seven datasets and 82 distinct test scenarios—demonstrates MLMP’s broad applicability and establishes a rigorous evaluation protocol for future work in adaptive, language-aware segmentation. While MLMP demonstrates strong, consistent gains, there remain opportunities to further refine its components. In particular, our current layer-weighting mechanism relies on entropy estimates from a shared projection head, which may not fully reflect each layer’s unique characteristics. Future work could investigate more flexible architectures—such as lightweight adapters or dedicated projection modules per layer—to more accurately assess and fuse intermediate features.
Appendix A Proof of Proposition 1
Unbiasedness and Variance Bound Proposition 1 (Restated): Assume each per-template gradient $g_{t}(\theta)=\nabla_{\theta}[L_{\text{UAML}}(T_{t})+L_{\text{ILE}}(T_{t})]$ has variance bounded by $\sigma^{2}$. Then, the ensemble gradient, defined by $\nabla_{\theta}\mathcal{L}_{\mathrm{final}}=\frac{1}{T}\sum_{t=1}^{T}g_{t}(\theta)$, is unbiased and satisfies the following variance bound:
$\mathbb{E}[\nabla_{\theta}L_{\text{final}}]=\mathbb{E}[g_{t}(\theta)],\quad\text{Var}(\nabla_{\theta}L_{\text{final}})=\frac{1}{T^{2}}\sum_{t=1}^{T}\text{Var}(g_{t}(\theta))\leq\frac{\sigma^{2}}{T}.$
Step 1: Unbiasedness. By linearity of expectation, we have:
$\mathbb{E}\left[\nabla_{\theta}L_{\text{final}}\right]=\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}g_{t}(\theta)\right]=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[g_{t}(\theta)]=\mathbb{E}[g_{t}(\theta)].$
Hence, averaging the gradients from all $T$ templates gives an unbiased estimate of the true gradient of the final loss.
Step 2: Variance Bound. Assuming independence with $\operatorname{Var}(g_{t})\leq\sigma^{2}$, we get
$\text{Var}(\nabla_{\theta}L_{\text{final}})=\text{Var}\left(\frac{1}{T}\sum_{t=1}^{T}g_{t}(\theta)\right)=\frac{1}{T^{2}}\sum_{t=1}^{T}\text{Var}(g_{t}(\theta))\leq\frac{T\sigma^{2}}{T^{2}}=\frac{\sigma^{2}}{T}$
Thus, the variance bound holds:
$\text{Var}(\nabla_{\theta}L_{\text{final}})\leq\frac{\sigma^{2}}{T}.$
This completes the proof of Proposition 1.
Appendix B Implementation Details
Unless otherwise noted, all experiments utilize NACLIP [30] with ViT-L/14 as the OVSS backbone. We adapt only the LayerNorm parameters within the vision encoder, amounting to approximately 0.02% of the model’s total parameters. Our adaptation setup follows prior work in classification [5, 31], using the Adam optimizer with a fixed learning rate of $0.001$ and 10 adaptation steps (iterations) across all experiments. Additionally we use batch of 2 images during adaptation. After each batch, model weights are explicitly reset to their original pre-adaptation values to ensure that each batch is adapted independently, without leveraging information from previously processed data. Following standard settings from  [1], we use the default softmax temperature value of $100$ in all experiments. All images are resized to $224\times 224$ pixels. Due to the high resolution of images in the Cityscapes dataset, we split them into overlapping patches of size $224\times 224$ pixels with an overlap of 112 pixels between patches. The segmentation predictions from these patches are aggregated to reconstruct the final, full-resolution segmentation maps. No image augmentation techniques are applied during either the adaptation or evaluation phases.
All experiments are conducted on NVIDIA V100 GPUs equipped with 32GB memory. We implement our approach using the PyTorch deep learning framework. To ensure statistical robustness and fairness in our comparisons, we repeat each experiment three times, reporting average performance along with standard deviation. We have provided detailed instructions and step-by-step scripts in our repository, clearly demonstrating how to generate datasets, perform the described test-time adaptations, and reproduce our reported results.
We adapt other baseline methods (TENT [8], TPT [6], WATT [5], CLIPArTT [31]) to work with spatial tokens. General adaptation hyperparameters (optimizer, learning rate, adaptation steps, and batch size) remain consistent with our setup. Method-specific hyperparameters or components are retained as reported in their original implementations. Specifically, for WATT, we used the sequential version (WATT-S) with default values of $l=2$ and $m=5$, for CLIPArTT we used the default $k=3$, and for TPT we used the 4 learnable tokens. Adapting these baseline methods to the segmentation task allowed us to systematically evaluate how various adaptation strategies—such as prompt tuning (TPT), pseudo-labeling (WATT, CLIPArTT), prompt refinement (CLIPArTT), and weight averaging (WATT)—perform in the context of open-vocabulary segmentation.
Appendix C Template Details
Table 4 lists the seven prompt templates used in our MLMP method. These templates were selected by the original CLIP authors [^1] https://github.com/openai/CLIP/blob/main/notebooks/Interacting_with_CLIP.ipynb and are general-purpose, not tailored to any specific image content. The CLIP repository also provides a full set of 80 prompt templates, which we use for larger-scale ablations.
More specifically, for the ablation in Fig. 4 of the main paper, we vary the number of templates $T$ as follows:
  • $T=1$: the default CLIP prompt “a photo of a {class}
  • $T=3$: the first three templates $\{T^{1},T^{2},T^{3}\}$ in Table 4
  • $T=7$: all seven templates $\{T^{1},\dots,T^{7}\}$ in Table 4
  • $T=20$: the first 20 templates from the 80-template pool
  • $T=80$: the complete set of 80 CLIP templates
Table 4: List of the seven prompt templates used in our MLMP method. These general-purpose templates, originally proposed by the CLIP authors, serve as diverse textual views of each class and are not tailored to specific datasets or domains.
ID Prompt template
$T^{1}$ itap of a {class}
$T^{2}$ a bad photo of the {class}.
$T^{3}$ a origami {class}.
$T^{4}$ a photo of the large {class}.
$T^{5}$ a {class} in a video game.
$T^{6}$ art of the {class}.
$T^{7}$ a photo of the small {class}.
Appendix D Dataset Details
This section details the datasets used in our experiments, along with the synthetic corruptions applied to evaluate robustness under distribution shifts.
More specifically, we conduct experiments on a diverse set of segmentation benchmarks:
  • Pascal VOC (V20/V21) [34]: Contains 20 foreground object classes (v20) with a background class (v21), widely used for benchmarking semantic segmentation tasks.
  • Pascal Context (P59/P60) [35]: An extension of the Pascal VOC 2010 dataset, providing pixel-level annotations for more than 400 classes. Due to sparsity, a frequently used subset includes 59 object classes plus a background class, totaling 60 categories.
  • Cityscapes [36]: A large-scale dataset for semantic segmentation of urban street scenes. It comprises around 5,000 finely annotated images from 50 cities, recorded under various daylight conditions, featuring dynamic objects, varying layouts, and changing backgrounds, capturing significant natural domain shifts.
  • COCO-Stuff [37]: Extends the original COCO dataset by adding annotations for background categories ("stuff"), resulting in 171 classes—80 objects, and 91 stuff categories.
  • COCO-Object [38]: A subset of the original COCO dataset, consisting exclusively of 80 object categories without background annotations.
In addition to the original versions of each dataset—some of which already reflect natural domain shifts (e.g., Cityscapes)—we further assess the robustness of all methods by applying synthetic corruptions. Inspired by ImageNet-C [33], we apply 15 synthetic corruptions to evaluate the robustness of segmentation models under various perturbations. These corruptions include Gaussian noise, shot noise, impulse noise, defocus blur, glass blur, motion blur, zoom blur, snow, frost, fog, brightness variations, contrast variations, elastic transformations, pixelation, and JPEG compression. Each corruption is applied at severity level 5, representing the most challenging scenario.
We resize images from all datasets to $224\times 224$ pixels. Due to the high resolution of Cityscapes images, we process them as overlapping patches of size $224\times 224$ pixels (with overlaps of 112 pixels). Predictions for these patches are subsequently aggregated to produce the final segmentation maps.
This curated set of 82 distinct test scenarios rigorously benchmarks our method across a comprehensive spectrum of conditions. It encompasses not only typical natural domain shifts and synthetic corruptions but also significant variations in image resolution, scene complexity, object scale, and semantic granularity.
Appendix E Computational Complexity
In this section, we provide an exhaustive analysis of each test-time adaptation method’s resource footprint by measuring: (i) latency, (ii) floating-point operations (FLOPs), (iii) peak GPU memory usage, and (iv) number of learnable parameters. For a fair comparison, all measurements were performed on a single test sample using the same NVIDIA V100 (32 GB) GPU. The results are summarized in Table 5.
We compare TENT, TPT, CLIPArTT, WATT, and our MLMP across all four complexity metrics. In TENT, WATT, CLIPArTT, and MLMP, only the LayerNorm parameters of the vision encoder are updated during adaptation, whereas TPT introduces additional learnable tokens at the input of the text encoder. Additionally TENT, WATT, and MLMP use a fixed sets of text features without a need for recalculating them during adaptation/evaluation, so all prompt templates can be encoded once (✓ in the “One-time Text Encoder” column) and then reused throughout both adaptation and evaluation phases. In contrast, TPT and CLIPArTT modify prompt embeddings or refine text templates during adaptation, requiring multiple forward passes through the text encoder.
In terms of latency, MLMP completes both adaptation and evaluation in just 0.582 ms, which is only marginally slower than TENT (0.480 ms) but substantially faster than WATT (5.215 ms) and CLIPArTT (3.525 ms). Despite leveraging multi-level fusion and multiple prompt templates, MLMP maintains a lightweight computational profile with only 82.4 GFLOPs and 82.9 GFLOPs for adaptation and evaluation, respectively—comparable to TENT and significantly lower than all other baselines. Furthermore, MLMP’s peak memory usage is among the lowest at 2,093.9 MB, and like TENT, it updates only 0.02% of the model parameters. These results highlight MLMP’s efficiency: it delivers rich representational capacity through multi-level and multi-prompt integration by boosting the results significantly (as shown in Table 4 of the main paper), yet remains almost as lightweight as the simplest baseline.
Table 5: Computational-complexity comparison across methods. For GFLOPs, only the forward-pass cost in adaptation and evaluation is measured; by common practice, the cost of back-propagation in adaptation phase can be approximated as twice the forward cost. A ✓ in the second column indicates that all text information are encoded once and cached. A dagger ($\dagger$) indicates that TPT adds additional parameters beyond the original network.
Method One-time Text Encoder Time (sec.) $\downarrow$ GFLOPs $\downarrow$ Max Memory (MB) $\downarrow$ Learn. Params (Ratio) $\downarrow$
Adapt Eval Adapt Eval
TENT 0.462 0.018 79.1 79.1 2,068.4 102,400 (0.02%)
TPT 0.445 0.031 275.6 275.6 2,583.1 3,072 (<0.01%)
CLIPArTT 3.494 0.031 1,755.5 275.6 8,928.5 102,400 (0.02%)
WATT 5.197 0.018 553.9 79.1 7,232.4 716,800 (0.17%)
MLMP (ours) 0.541 0.041 82.4 82.9 2,093.9 102,400 (0.02%)
Appendix F Performance with a Single Test Sample
As shown in Table 6, while TENT yields improvements on the original dataset, it leads to a performance drop on V20-C. In contrast, our method, MLMP, consistently improves performance over the no adaptation baseline, with gains of 8.77% on the otiginal data and 9.40% on the average across corruptions.
Table 6: mIoU performance comparison when using a single test sample for V20 dataset.
OVSS Backbone: NACLIP Adaptation Method
Dataset: V20 No Adapt. TENT MLMP
Original 75.91 76.20 84.68
Gaussian noise 62.89 62.59 71.92
Shot noise 66.26 65.45 75.78
Impulse noise 63.16 63.34 72.35
Defocus blur 72.59 71.37 80.91
Glass blur 71.44 69.95 79.29
Motion blur 73.10 72.55 81.64
Zoom blur 59.03 60.64 69.99
Snow 71.49 70.03 80.64
Frost 65.38 65.96 74.33
Fog 70.69 69.39 80.77
Brightness 74.95 74.73 84.58
Contrast 71.51 69.74 79.78
Elastic transform 62.86 65.09 75.32
Pixelate 77.28 75.84 85.63
JPEG compression 72.59 70.85 83.20
V20-C Average 69.01 68.50 78.41
Appendix G Generalization Across Model Variants
To evaluate the robustness and generality of our method, we conduct a series of experiments using different model configurations within the segmentation pipeline. Specifically, we assess how MLMP performs when (i) changing the vision transformer backbone, (ii) switching between different open-vocabulary semantic segmentation (OVSS) formulations, and (iii) adopting an entirely different vision–language model. These experiments demonstrate that our method maintains consistent improvements across a wide range of architectural and algorithmic configurations, confirming its flexibility and transferability.
G.1 Comparison Across ViT Backbones
To evaluate whether our method generalizes across different vision transformer backbones, we replicate the main experiments using ViT-B/16 and ViT-B/32 in place of the default ViT-L/14 model. These backbones represent lighter configurations, with fewer parameters and larger patch (32 and 16). As shown in Tables 7 and 8, MLMP continues to provide substantial improvements over all baselines across both configurations, on both clean data (V20 Original) and under severe synthetic corruptions. These results confirm that the benefits of our multi-level and multi-prompt adaptation strategy are not tied to model scale or specific architectural configurations, and remain effective even in lower-resolution settings.
Table 7: Performance comparison of test-time adaptation methods using the ViT-B/16 backbone with NaCLIP as the OVSS model. Results are reported as mIoU scores on the V20 dataset (original) and 15 corruption types. MLMP achieves the highest performance across all settings, demonstrating strong robustness even with a smaller backbone.
OVSS Backbone: NACLIP Adaptation Method
Dataset: V20 No Adapt. TENT TPT WATT CLIPArTT MLMP
Original 77.62 79.15 $\pm$0.06 77.63 $\pm$0.01 43.18 $\pm$0.07 72.70 $\pm$0.17 84.18 $\pm$0.07
Gaussian noise 48.00 52.04 $\pm$0.12 48.11 $\pm$0.00 18.83 $\pm$0.14 33.80 $\pm$0.52 61.67 $\pm$0.00
Shot noise 52.49 55.56 $\pm$0.16 52.41 $\pm$0.00 19.86 $\pm$0.12 37.62 $\pm$0.33 64.97 $\pm$0.10
Impulse noise 49.51 52.87 $\pm$0.11 49.41 $\pm$0.01 19.12 $\pm$0.23 35.92 $\pm$0.05 61.15 $\pm$0.09
Defocus blur 68.03 69.85 $\pm$0.04 67.88 $\pm$0.01 37.54 $\pm$0.17 56.77 $\pm$0.19 76.71 $\pm$0.02
Glass blur 62.17 65.14 $\pm$0.17 62.45 $\pm$0.00 31.12 $\pm$0.08 47.86 $\pm$0.15 72.62 $\pm$0.04
Motion blur 69.56 71.93 $\pm$0.02 69.54 $\pm$0.01 36.89 $\pm$0.15 58.64 $\pm$0.29 77.08 $\pm$0.03
Zoom blur 47.34 52.30 $\pm$0.19 47.38 $\pm$0.01 22.83 $\pm$0.12 33.52 $\pm$0.20 59.05 $\pm$0.20
Snow 60.88 64.38 $\pm$0.05 61.24 $\pm$0.00 27.68 $\pm$0.21 49.91 $\pm$0.38 71.41 $\pm$0.15
Frost 55.45 58.38 $\pm$0.14 55.44 $\pm$0.00 29.66 $\pm$0.23 46.94 $\pm$0.04 67.42 $\pm$0.09
Fog 67.07 70.01 $\pm$0.02 67.07 $\pm$0.00 35.84 $\pm$0.23 59.98 $\pm$0.09 76.32 $\pm$0.02
Brightness 73.33 75.14 $\pm$0.17 73.23 $\pm$0.00 40.87 $\pm$0.09 67.30 $\pm$0.33 82.14 $\pm$0.03
Contrast 60.30 63.30 $\pm$0.04 60.20 $\pm$0.00 29.68 $\pm$0.12 48.42 $\pm$0.55 70.02 $\pm$0.04
Elastic transform 50.14 54.83 $\pm$0.01 50.00 $\pm$0.01 23.32 $\pm$0.14 43.67 $\pm$0.00 63.65 $\pm$0.07
Pixelate 75.48 77.44 $\pm$0.03 75.31 $\pm$0.01 42.26 $\pm$0.08 67.33 $\pm$0.08 83.29 $\pm$0.01
JPEG compression 69.17 70.97 $\pm$0.07 69.15 $\pm$0.01 34.94 $\pm$0.07 60.30 $\pm$0.54 79.35 $\pm$0.12
V20-C Average 60.59 63.61 60.59 30.03 49.87 71.12
Table 8: Performance comparison of test-time adaptation methods using the ViT-B/32 backbone with NaCLIP as the OVSS model. Results are reported as mIoU scores on the V20 dataset (original) and 15 corruption types. MLMP achieves the highest performance across all settings, demonstrating strong robustness even with a smaller backbone and smaller patch size.
OVSS Backbone: NACLIP Adaptation Method
Dataset: V20 No Adapt. TENT TPT WATT CLIPArTT MLMP
Original 72.43 72.83  $\pm$0.01 72.52  $\pm$0.00 50.71  $\pm$0.10 67.67  $\pm$0.15 79.95  $\pm$0.01
Gaussian noise 47.59 49.30  $\pm$0.18 47.43  $\pm$0.00 29.26  $\pm$0.09 37.36  $\pm$0.17 59.27  $\pm$0.16
Shot noise 51.80 54.43  $\pm$0.21 51.83  $\pm$0.00 31.82  $\pm$0.15 41.07  $\pm$0.21 63.73  $\pm$0.12
Impulse noise 48.79 51.59  $\pm$0.11 48.79  $\pm$0.00 30.31  $\pm$0.04 37.65  $\pm$0.17 58.81  $\pm$0.04
Defocus blur 60.23 61.86  $\pm$0.07 60.17  $\pm$0.00 36.54  $\pm$0.20 48.97  $\pm$0.21 66.70  $\pm$0.08
Glass blur 54.59 56.73  $\pm$0.09 54.80  $\pm$0.00 28.70  $\pm$0.10 37.96  $\pm$0.26 65.34  $\pm$0.03
Motion blur 59.53 60.49  $\pm$0.02 59.65  $\pm$0.00 35.80  $\pm$0.06 47.61  $\pm$0.52 66.71  $\pm$0.08
Zoom blur 38.66 41.07  $\pm$0.10 38.62  $\pm$0.00 21.72  $\pm$0.14 24.39  $\pm$0.18 49.53  $\pm$0.08
Snow 49.17 51.01  $\pm$0.08 48.88  $\pm$0.01 27.55  $\pm$0.17 40.30  $\pm$0.00 62.30  $\pm$0.09
Frost 47.60 50.18  $\pm$0.09 47.53  $\pm$0.00 28.48  $\pm$0.15 39.23  $\pm$0.12 60.22  $\pm$0.07
Fog 56.25 59.47  $\pm$0.08 56.22  $\pm$0.00 35.76  $\pm$0.09 47.63  $\pm$0.29 68.18  $\pm$0.04
Brightness 68.03 68.94  $\pm$0.04 67.95  $\pm$0.00 46.11  $\pm$0.02 61.07  $\pm$0.25 76.90  $\pm$0.09
Contrast 48.79 50.67  $\pm$0.07 48.88  $\pm$0.02 29.54  $\pm$0.06 36.64  $\pm$0.05 58.81  $\pm$0.11
Elastic transform 52.73 55.59  $\pm$0.20 52.75  $\pm$0.00 29.86  $\pm$0.21 44.84  $\pm$0.18 65.39  $\pm$0.09
Pixelate 68.61 69.15  $\pm$0.01 68.65  $\pm$0.01 44.57  $\pm$0.17 60.30  $\pm$0.41 77.28  $\pm$0.04
JPEG compression 63.86 65.12  $\pm$0.05 63.87  $\pm$0.00 42.85  $\pm$0.11 55.44  $\pm$0.08 74.40  $\pm$0.09
V20-C Average 54.42 56.37 54.40 33.26 44.03 64.97
G.2 Comparison Across OVSS Methods
Our method is designed to be flexible and agnostic to the underlying open-vocabulary semantic segmentation (OVSS) formulation. While all main experiments in the paper use NaCLIP as the OVSS baseline—paired with No Adapt., TENT, TPT, CLIPArTT, WATT, and our MLMP—we further evaluate the generality of our approach by applying MLMP to two alternative OVSS methods.
Specifically, we consider the original CLIP [1], adapted for pixel-wise segmentation via patch-level similarity, and SCLIP [27], which incorporates spatial priors into the vision-language matching process. As shown in Table 9, applying MLMP on top of these OVSS baselines yields consistent improvements across both clean and corrupted settings.
Table 9: Performance comparison of test-time adaptation methods using the original CLIP [1] and SCLIP [27] as the OVSS model with a ViT-B/16 backbone. MLMP consistently improves performance across all settings, highlighting its strong generalization capabilities.
OVSS: CLIP [1] Adaptation Method
Dataset: V20 No Adapt. TENT MLMP
Original 33.11 51.36  $\pm$0.00 61.47  $\pm$0.01
Gaussian noise 22.74 37.78  $\pm$0.11 49.35  $\pm$0.02
Shot noise 23.67 38.86  $\pm$0.20 50.81  $\pm$0.05
Impulse noise 22.36 36.63  $\pm$0.15 47.81  $\pm$0.17
Defocus blur 31.83 48.33  $\pm$0.13 58.25  $\pm$0.04
Glass blur 28.60 46.59  $\pm$0.29 56.02  $\pm$0.02
Motion blur 32.62 50.61  $\pm$0.07 59.55  $\pm$0.31
Zoom blur 23.42 39.43  $\pm$0.06 47.65  $\pm$0.10
Snow 28.38 48.05  $\pm$0.04 55.89  $\pm$0.04
Frost 26.20 45.33  $\pm$0.04 53.56  $\pm$0.17
Fog 28.66 46.98  $\pm$0.05 56.62  $\pm$0.02
Brightness 33.71 51.77  $\pm$0.24 60.96  $\pm$0.08
Contrast 26.06 42.86  $\pm$0.08 53.55  $\pm$0.04
Elastic transform 27.12 45.92  $\pm$0.11 51.02  $\pm$0.21
Pixelate 33.32 51.11  $\pm$0.00 61.22  $\pm$0.09
JPEG compression 31.64 49.76  $\pm$0.04 59.79  $\pm$0.06
V21-C Average 28.02 45.33 54.80
OVSS: SCLIP [27] Adaptation Method
Dataset: V20 No Adapt. TENT MLMP
Original 78.20 79.12  $\pm$0.05 84.91  $\pm$0.01
Gaussian noise 45.65 49.72  $\pm$0.07 61.20  $\pm$0.09
Shot noise 50.21 54.09  $\pm$0.15 64.06  $\pm$0.09
Impulse noise 47.05 50.62  $\pm$0.05 59.73  $\pm$0.14
Defocus blur 66.40 66.44  $\pm$0.05 75.03  $\pm$0.07
Glass blur 62.01 64.79  $\pm$0.22 72.22  $\pm$0.27
Motion blur 68.95 70.81  $\pm$0.02 76.23  $\pm$0.10
Zoom blur 45.12 48.50  $\pm$0.11 57.84  $\pm$0.16
Snow 60.61 64.60  $\pm$0.20 71.70  $\pm$0.06
Frost 56.14 58.43  $\pm$0.03 68.57  $\pm$0.02
Fog 68.34 70.03  $\pm$0.17 76.57  $\pm$0.06
Brightness 74.03 74.62  $\pm$0.11 82.91  $\pm$0.04
Contrast 58.98 61.69  $\pm$0.08 69.57  $\pm$0.10
Elastic transform 52.29 56.45  $\pm$0.02 65.05  $\pm$0.14
Pixelate 75.56 76.66  $\pm$0.09 82.99  $\pm$0.00
JPEG compression 68.38 69.85  $\pm$0.12 78.69  $\pm$0.17
V21-C Average 59.98 62.49 70.82
G.3 Generalization to Emerging VLMs
To further assess the generality of our approach, we evaluate MLMP on SigLIP v2 [39], one of the most recently introduced vision–language models. As shown in Table 10, MLMP continues to deliver consistent improvements across both natural and corrupted datasets, indicating that the core components of our method—multi-level and multi-prompt aggregation—generalize well beyond CLIP-based architectures.
Table 10: Performance comparison of test-time adaptation using SigLIP-2 as the OVSS model. Results are reported as mIoU scores on the V20 dataset. MLMP shows improvement over the baseline in most corruption types.
OVSS Backbone: SigLIP-2 Adaptation Method
Dataset: V20 No Adapt. MLMP
Original 66.62 67.88  $\pm$0.32
Gaussian noise 39.74 41.05  $\pm$0.34
Shot noise 40.24 43.56  $\pm$0.07
Impulse noise 40.76 42.43  $\pm$0.22
Defocus blur 47.15 50.41  $\pm$0.02
Glass blur 38.64 40.60  $\pm$0.55
Motion blur 42.30 42.98  $\pm$0.39
Zoom blur 27.29 29.20  $\pm$0.03
Snow 03.85 04.84  $\pm$0.03
Frost 21.32 23.12  $\pm$0.02
Fog 70.73 70.86  $\pm$0.04
Brightness 18.45 19.14  $\pm$0.27
Contrast 70.88 70.35  $\pm$0.06
Elastic transform 38.41 38.71  $\pm$0.05
Pixelate 57.85 59.14  $\pm$0.07
JPEG compression 56.35 58.55  $\pm$0.04
V20-C Average 40.93 42.33
Appendix H Visualization of Layer-Wise Confidence Weights
In the main paper (Fig. 3), we presented the layer-wise confidence weights for the V20, Cityscapes, and COCO-Stuff datasets. Here, we extend the analysis by visualizing the weights for four additional datasets: V21, P59, P60, and COCO-Object in Fig. 6. As with the earlier results, we plot the mean and standard deviation of the learned weights across layers under various corruption types.
The observed trends closely mirror those reported in the main paper. In particular, deeper layers consistently receive higher confidence scores, while lower and mid-level layers still contribute under corrupted conditions. This reweighting behavior reflects the adaptive nature of our fusion strategy, which dynamically emphasizes the most reliable representations depending on the dataset and corruption type. These results further support the robustness and generality of our layer-wise uncertainty-aware fusion mechanism across diverse segmentation benchmarks.
Figure 6: Mean and standard deviation of weights of intermediate layers for several datasets.
Appendix I Visualization of Segmentation Maps
Figure 7 presents qualitative segmentation results on the v20 dataset, comparing predictions from the non-adaptive baseline, TENT, and our MLMP method. MLMP demonstrates stronger spatial consistency and fewer semantic errors, particularly in challenging regions where both the baseline and TENT struggle. The integration of intermediate-layer supervision appears especially beneficial for refining small object boundaries and correcting fine-grained details.
In Figures 816, we provide additional qualitative examples across a range of corruption types, using NaCLIP as the our OVSS backbone. These include both successful and failure cases under Gaussian noise, defocus blur, snow, and JPEG compression. Overall, the results illustrate the robustness of MLMP in mitigating noise-induced artifacts and improving prediction confidence, while also highlighting failure modes that warrant further investigation.
Figure 7: Qualitative results for No Adapt., TENT and MLMP on V20 Original.
Figure 8: Failed Cases of MLMP on V20 Original.
Figure 9: Good Cases of MLMP on V20 for V20 gaussian noise corruption.
Figure 10: Failed Cases of MLMP on V20 gaussian noise corruption.
Figure 11: Good Cases of MLMP on V20 defocus blur corruption.
Figure 12: Failed Cases of MLMP on V20 defocus blur corruption.
Figure 13: Good Cases of MLMP on V20 snow corruption.
Figure 14: Failed Cases of MLMP on V20 snow corruption.
Figure 15: Good Cases of MLMP on V20 JPEG compression corruption.
Figure 16: Failed Cases of MLMP on V20 JPEG compression corruption.
Appendix J Detailed Results of the Main Paper
This section provides complete versions of the experimental results that were summarized or partially reported in the main paper. It includes full tables for ablation studies and detailed comparisons with baseline adaptation methods across various datasets and evaluation settings. These results offer a more comprehensive view of the effectiveness and robustness of our proposed approach.
J.1 Ablation Studies
We provide here the detailed versions of the tables referenced in the ablation section of the main paper. Table 11 shows the detailed results of using different layer ranges in the proposed multi-level adaptation. Table 12 reports results for different strategies to integrate different prompt templates. Table 13 presents a detailed analysis of the impact of the number of templates. Finally, Table 14 presents results for the combination of all proposed components, showing individual and combined contributions.
Table 11: mIoU performance over different layer aggregation strategies. Maximum value in each row is highlighted.
ViT-L/14 Layer Range $L_{24}$ (last) $L_{23\text{--}24}$ (last two) $L_{22\text{--}24}$ (last three) $L_{19\text{--}24}$ (last 25%) $L_{13\text{--}24}$ (last 50%) $L_{7\text{--}24}$ (last 75%) $L_{1\text{--}24}$ (all layers)
Original (V20) 77.00 $\pm$0.04 77.65 $\pm$0.02 77.66 $\pm$0.09 80.61 $\pm$0.05 80.50 $\pm$0.03 81.67 $\pm$0.04 78.79 $\pm$0.02
Gaussian noise 63.02 $\pm$0.06 64.41 $\pm$0.05 65.39 $\pm$0.13 66.88 $\pm$0.18 66.88 $\pm$0.02 67.82 $\pm$0.01 63.06 $\pm$0.09
Shot noise 65.88 $\pm$0.06 67.74 $\pm$0.11 68.43 $\pm$0.04 70.11 $\pm$0.09 70.40 $\pm$0.02 70.52 $\pm$0.12 64.88 $\pm$0.02
Impulse noise 64.17 $\pm$0.04 65.53 $\pm$0.09 66.19 $\pm$0.12 67.24 $\pm$0.17 67.15 $\pm$0.09 68.10 $\pm$0.01 62.09 $\pm$0.05
Defocus blur 72.06 $\pm$0.12 72.93 $\pm$0.19 72.84 $\pm$0.02 76.10 $\pm$0.16 76.37 $\pm$0.05 78.78 $\pm$0.02 77.56 $\pm$0.09
Glass blur 70.74 $\pm$0.07 71.73 $\pm$0.06 72.53 $\pm$0.15 74.20 $\pm$0.12 75.53 $\pm$0.05 77.66 $\pm$0.05 75.85 $\pm$0.02
Motion blur 73.50 $\pm$0.10 73.74 $\pm$0.08 74.62 $\pm$0.09 76.83 $\pm$0.07 77.50 $\pm$0.07 79.39 $\pm$0.16 77.22 $\pm$0.15
Zoom blur 61.36 $\pm$0.07 62.10 $\pm$0.01 62.56 $\pm$0.00 63.99 $\pm$0.20 64.59 $\pm$0.14 66.50 $\pm$0.16 64.79 $\pm$0.01
Snow 71.04 $\pm$0.05 72.09 $\pm$0.04 72.56 $\pm$0.02 74.47 $\pm$0.12 74.41 $\pm$0.01 76.39 $\pm$0.02 73.72 $\pm$0.07
Frost 67.01 $\pm$0.02 66.92 $\pm$0.02 67.23 $\pm$0.04 68.94 $\pm$0.01 70.03 $\pm$0.01 71.17 $\pm$0.00 68.23 $\pm$0.02
Fog 70.54 $\pm$0.07 71.44 $\pm$0.07 71.50 $\pm$0.05 74.50 $\pm$0.16 74.62 $\pm$0.08 76.55 $\pm$0.06 74.37 $\pm$0.03
Brightness 75.61 $\pm$0.02 76.27 $\pm$0.00 76.80 $\pm$0.04 79.16 $\pm$0.07 79.64 $\pm$0.16 81.34 $\pm$0.04 79.11 $\pm$0.05
Contrast 70.51 $\pm$0.04 71.72 $\pm$0.13 72.17 $\pm$0.00 74.58 $\pm$0.08 75.42 $\pm$0.09 76.87 $\pm$0.06 74.09 $\pm$0.00
Elastic transform 65.78 $\pm$0.05 66.08 $\pm$0.08 66.06 $\pm$0.06 68.62 $\pm$0.15 70.38 $\pm$0.09 70.59 $\pm$0.14 67.91 $\pm$0.12
Pixelate 76.95 $\pm$0.12 78.38 $\pm$0.02 78.46 $\pm$0.05 80.70 $\pm$0.05 81.11 $\pm$0.03 83.02 $\pm$0.04 81.53 $\pm$0.02
JPEG compression 71.84 $\pm$0.15 73.88 $\pm$0.11 74.40 $\pm$0.07 76.96 $\pm$0.02 77.67 $\pm$0.03 78.73 $\pm$0.08 75.87 $\pm$0.19
V20-C Average 69.33 70.33 70.78 72.89 73.45 74.90 72.02
Table 12: mIoU performance comparison for different strategies to integrate different prompt templates.
Dataset: V20 Text Params Loss
Original 78.91 $\pm$0.07 74.46 $\pm$0.21 79.70 $\pm$0.06
Gaussian noise 66.27 $\pm$0.00 62.83 $\pm$0.04 66.75 $\pm$0.01
Shot noise 69.78 $\pm$0.10 67.10 $\pm$0.12 70.03 $\pm$0.04
Impulse noise 66.73 $\pm$0.03 64.57 $\pm$0.52 67.88 $\pm$0.07
Defocus blur 74.05 $\pm$0.10 70.28 $\pm$0.16 74.31 $\pm$0.09
Glass blur 73.31 $\pm$0.08 70.28 $\pm$0.34 74.38 $\pm$0.36
Motion blur 75.20 $\pm$0.09 71.56 $\pm$0.09 75.72 $\pm$0.04
Zoom blur 63.31 $\pm$0.11 61.29 $\pm$0.09 64.61 $\pm$0.01
Snow 73.78 $\pm$0.02 70.10 $\pm$0.30 74.66 $\pm$0.01
Frost 68.90 $\pm$0.06 66.42 $\pm$0.04 69.50 $\pm$0.01
Fog 72.60 $\pm$0.03 67.63 $\pm$0.07 73.33 $\pm$0.03
Brightness 78.16 $\pm$0.02 74.05 $\pm$0.03 78.69 $\pm$0.02
Contrast 73.75 $\pm$0.06 69.47 $\pm$0.24 74.09 $\pm$0.01
Elastic transform 68.41 $\pm$0.02 64.98 $\pm$0.09 69.14 $\pm$0.05
Pixelate 79.59 $\pm$0.06 75.54 $\pm$0.05 80.09 $\pm$0.03
JPEG compression 74.98 $\pm$0.05 70.55 $\pm$0.11 75.56 $\pm$0.02
V20-C Average 71.92 68.44 72.58
Table 13: IoU performance of our method for different numbers of templates.
Dataset: V20 1 Template 3 Templates 7 Templates 20 Templates 80 Templates
Original 77.00 $\pm$0.04 79.48 $\pm$0.08 79.70 $\pm$0.06 79.17 $\pm$0.01 79.25 $\pm$0.02
Gaussian noise 63.02 $\pm$0.06 66.51 $\pm$0.10 66.75 $\pm$0.01 65.94 $\pm$0.09 66.02 $\pm$0.02
Shot noise 65.88 $\pm$0.06 70.20 $\pm$0.03 70.03 $\pm$0.04 68.59 $\pm$0.01 69.00 $\pm$0.01
Impulse noise 64.17 $\pm$0.04 67.57 $\pm$0.01 67.88 $\pm$0.07 66.53 $\pm$0.06 66.91 $\pm$0.02
Defocus blur 72.06 $\pm$0.12 74.66 $\pm$0.02 74.31 $\pm$0.09 73.89 $\pm$0.12 74.09 $\pm$0.06
Glass blur 70.74 $\pm$0.07 74.38 $\pm$0.12 74.38 $\pm$0.36 73.78 $\pm$0.09 73.88 $\pm$0.07
Motion blur 73.50 $\pm$0.10 75.80 $\pm$0.08 75.72 $\pm$0.04 75.19 $\pm$0.04 75.19 $\pm$0.04
Zoom blur 61.36 $\pm$0.07 63.47 $\pm$0.02 64.61 $\pm$0.01 63.50 $\pm$0.03 63.84 $\pm$0.04
Snow 71.04 $\pm$0.05 74.09 $\pm$0.08 74.66 $\pm$0.01 73.78 $\pm$0.03 73.91 $\pm$0.01
Frost 67.01 $\pm$0.02 69.09 $\pm$0.01 69.50 $\pm$0.01 69.08 $\pm$0.00 69.10 $\pm$0.02
Fog 70.54 $\pm$0.07 73.69 $\pm$0.01 73.33 $\pm$0.03 72.87 $\pm$0.03 72.86 $\pm$0.05
Brightness 75.61 $\pm$0.02 78.49 $\pm$0.05 78.69 $\pm$0.02 77.86 $\pm$0.02 78.06 $\pm$0.02
Contrast 70.51 $\pm$0.04 74.01 $\pm$0.22 74.09 $\pm$0.01 73.48 $\pm$0.04 73.56 $\pm$0.09
Elastic transform 65.78 $\pm$0.05 68.09 $\pm$0.01 69.14 $\pm$0.05 67.94 $\pm$0.02 68.31 $\pm$0.09
Pixelate 76.95 $\pm$0.12 79.91 $\pm$0.04 80.09 $\pm$0.03 79.15 $\pm$0.06 79.41 $\pm$0.05
JPEG compression 71.84 $\pm$0.15 75.00 $\pm$0.05 75.56 $\pm$0.02 74.45 $\pm$0.09 74.75 $\pm$0.01
V20-C Average 69.33 72.33 72.58 71.74 71.93
Table 14: Detailed mIoU comparison of MLMP components, showing individual and combined contributions.
Multi-Level Fusion
Multi-Prompt Loss
Image-Level Entropy
Uncertainty-Aware Weight.
Original 77.00 $\pm$0.04 77.38 $\pm$0.01 81.67 $\pm$0.04 79.70 $\pm$0.06 78.74 $\pm$0.08 78.97 $\pm$0.03 83.00 $\pm$0.03 77.69 $\pm$0.01 82.70 $\pm$0.01 81.15 $\pm$0.02 79.13 $\pm$0.00 83.76 $\pm$0.00
Gaussian noise 63.02 $\pm$0.06 65.42 $\pm$0.04 67.82 $\pm$0.01 66.75 $\pm$0.01 65.66 $\pm$0.04 65.96 $\pm$0.04 69.13 $\pm$0.07 66.17 $\pm$0.04 69.00 $\pm$0.05 69.62 $\pm$0.01 67.35 $\pm$0.09 71.13 $\pm$0.09
Shot noise 65.88 $\pm$0.06 67.49 $\pm$0.01 70.52 $\pm$0.12 70.03 $\pm$0.04 68.97 $\pm$0.03 69.05 $\pm$0.03 72.31 $\pm$0.01 69.50 $\pm$0.06 73.22 $\pm$0.03 72.89 $\pm$0.02 71.02 $\pm$0.05 75.02 $\pm$0.03
Impulse noise 64.17 $\pm$0.04 65.11 $\pm$0.17 68.10 $\pm$0.01 67.88 $\pm$0.07 66.35 $\pm$0.12 65.39 $\pm$0.12 68.86 $\pm$0.13 65.16 $\pm$0.10 68.77 $\pm$0.15 70.31 $\pm$0.09 67.17 $\pm$0.09 71.34 $\pm$0.11
Defocus blur 72.06 $\pm$0.12 76.65 $\pm$0.23 78.78 $\pm$0.02 74.31 $\pm$0.09 75.00 $\pm$0.07 76.46 $\pm$0.15 78.78 $\pm$0.10 77.29 $\pm$0.10 79.78 $\pm$0.05 77.14 $\pm$0.05 77.79 $\pm$0.19 80.36 $\pm$0.06
Glass blur 70.74 $\pm$0.07 75.24 $\pm$0.05 77.66 $\pm$0.05 74.38 $\pm$0.36 73.54 $\pm$0.23 75.46 $\pm$0.09 77.48 $\pm$0.01 75.71 $\pm$0.10 78.09 $\pm$0.04 76.17 $\pm$0.08 76.87 $\pm$0.07 78.84 $\pm$0.05
Motion blur 73.50 $\pm$0.10 77.16 $\pm$0.13 79.39 $\pm$0.16 75.72 $\pm$0.04 76.09 $\pm$0.09 77.72 $\pm$0.09 79.97 $\pm$0.04 78.81 $\pm$0.03 81.49 $\pm$0.02 78.25 $\pm$0.05 79.00 $\pm$0.07 81.41 $\pm$0.05
Zoom blur 61.36 $\pm$0.07 64.22 $\pm$0.12 66.50 $\pm$0.16 64.61 $\pm$0.01 64.04 $\pm$0.02 66.38 $\pm$0.15 68.41 $\pm$0.13 65.12 $\pm$0.16 67.69 $\pm$0.08 68.32 $\pm$0.17 67.61 $\pm$0.05 69.41 $\pm$0.12
Snow 71.04 $\pm$0.05 72.64 $\pm$0.07 76.39 $\pm$0.02 74.66 $\pm$0.01 74.16 $\pm$0.02 73.25 $\pm$0.12 77.31 $\pm$0.11 74.05 $\pm$0.08 78.50 $\pm$0.16 77.20 $\pm$0.02 74.94 $\pm$0.06 79.53 $\pm$0.05
Frost 67.01 $\pm$0.02 67.30 $\pm$0.03 71.17 $\pm$0.00 69.50 $\pm$0.01 69.31 $\pm$0.00 67.57 $\pm$0.19 71.73 $\pm$0.21 68.31 $\pm$0.12 72.81 $\pm$0.08 71.34 $\pm$0.02 68.69 $\pm$0.11 73.20 $\pm$0.07
Fog 70.54 $\pm$0.07 72.73 $\pm$0.03 76.55 $\pm$0.06 73.33 $\pm$0.03 73.41 $\pm$0.00 75.06 $\pm$0.08 78.38 $\pm$0.03 73.48 $\pm$0.13 77.62 $\pm$0.05 75.98 $\pm$0.02 75.81 $\pm$0.02 79.81 $\pm$0.06
Brightness 75.61 $\pm$0.02 77.23 $\pm$0.07 81.34 $\pm$0.04 78.69 $\pm$0.02 77.34 $\pm$0.01 77.69 $\pm$0.02 82.00 $\pm$0.03 77.20 $\pm$0.05 82.16 $\pm$0.03 80.63 $\pm$0.05 78.27 $\pm$0.06 83.51 $\pm$0.01
Contrast 70.51 $\pm$0.04 73.36 $\pm$0.08 76.87 $\pm$0.06 74.09 $\pm$0.01 74.14 $\pm$0.14 73.57 $\pm$0.05 77.42 $\pm$0.09 74.09 $\pm$0.11 78.09 $\pm$0.08 76.97 $\pm$0.01 74.75 $\pm$0.06 79.06 $\pm$0.16
Elastic transform 65.78 $\pm$0.05 65.38 $\pm$0.12 70.59 $\pm$0.14 69.14 $\pm$0.05 67.92 $\pm$0.06 68.76 $\pm$0.04 72.96 $\pm$0.02 66.78 $\pm$0.03 71.80 $\pm$0.02 71.53 $\pm$0.03 69.77 $\pm$0.07 74.03 $\pm$0.01
Pixelate 76.95 $\pm$0.12 79.53 $\pm$0.03 83.02 $\pm$0.04 80.09 $\pm$0.03 79.09 $\pm$0.03 80.77 $\pm$0.05 83.93 $\pm$0.06 79.85 $\pm$0.08 83.90 $\pm$0.00 81.92 $\pm$0.12 81.34 $\pm$0.07 84.97 $\pm$0.04
JPEG compression 71.84 $\pm$0.15 74.38 $\pm$0.12 78.73 $\pm$0.08 75.56 $\pm$0.02 74.77 $\pm$0.11 76.77 $\pm$0.00 80.81 $\pm$0.09 74.61 $\pm$0.06 79.79 $\pm$0.02 77.98 $\pm$0.02 77.94 $\pm$0.07 82.06 $\pm$0.01
V20-C Average 69.33 71.59 74.90 72.58 71.99 72.66 75.97 72.41 76.18 75.08 73.89 77.58
J.2 Comparison with Alternative Adaptation Methods
This section presents comprehensive tables with the full experimental results corresponding to those summarized in the main paper. Table 15 reports detailed results for the V21 dataset, Table 16 for P59, Table 17 for P60, and Table 18 for the Cityscapes dataset.
Table 15: mIoU comparison of MLMP and baseline methods on the V21 dataset, evaluated on both the original images and 15 corruption types.
OVSS Backbone: NACLIP Adaptation Method
Dataset: V21 No Adapt. TENT TPT WATT CLIPArTT MLMP
Original 45.12 45.65 $\pm$0.02 45.17 $\pm$0.00 28.58 $\pm$0.05 39.50 $\pm$0.04 50.78 $\pm$0.02
Gaussian noise 37.40 37.95 $\pm$0.00 37.34 $\pm$0.00 20.93 $\pm$0.07 30.05 $\pm$0.18 43.59 $\pm$0.01
Shot noise 39.33 39.17 $\pm$0.03 39.23 $\pm$0.00 22.06 $\pm$0.06 32.07 $\pm$0.17 45.55 $\pm$0.02
Impulse noise 37.81 37.73 $\pm$0.04 37.78 $\pm$0.00 19.95 $\pm$0.05 30.52 $\pm$0.08 43.89 $\pm$0.01
Defocus blur 41.46 41.46 $\pm$0.03 41.46 $\pm$0.00 24.79 $\pm$0.04 34.27 $\pm$0.05 46.00 $\pm$0.00
Glass blur 41.76 41.55 $\pm$0.01 41.82 $\pm$0.00 24.61 $\pm$0.03 34.55 $\pm$0.16 46.83 $\pm$0.04
Motion blur 42.65 42.81 $\pm$0.01 42.74 $\pm$0.00 25.66 $\pm$0.04 36.07 $\pm$0.11 47.72 $\pm$0.04
Zoom blur 34.46 34.25 $\pm$0.00 34.44 $\pm$0.00 20.74 $\pm$0.05 28.44 $\pm$0.07 39.07 $\pm$0.02
Snow 40.13 40.47 $\pm$0.00 40.23 $\pm$0.01 25.02 $\pm$0.06 33.98 $\pm$0.03 46.30 $\pm$0.05
Frost 40.70 41.83 $\pm$0.08 40.80 $\pm$0.00 23.43 $\pm$0.05 34.09 $\pm$0.01 46.78 $\pm$0.01
Fog 42.50 42.67 $\pm$0.00 42.47 $\pm$0.00 24.95 $\pm$0.05 36.37 $\pm$0.04 47.61 $\pm$0.06
Brightness 44.21 44.64 $\pm$0.03 44.27 $\pm$0.00 27.54 $\pm$0.07 38.44 $\pm$0.11 49.89 $\pm$0.08
Contrast 40.44 40.14 $\pm$0.03 40.44 $\pm$0.00 23.89 $\pm$0.04 33.68 $\pm$0.05 45.22 $\pm$0.00
Elastic transform 40.63 41.78 $\pm$0.01 40.67 $\pm$0.00 24.40 $\pm$0.05 35.38 $\pm$0.01 46.87 $\pm$0.02
Pixelate 44.70 44.95 $\pm$0.03 44.79 $\pm$0.00 27.74 $\pm$0.06 38.14 $\pm$0.06 50.11 $\pm$0.01
JPEG compression 43.05 42.87 $\pm$0.04 43.05 $\pm$0.00 26.04 $\pm$0.05 36.29 $\pm$0.10 48.27 $\pm$0.02
V21-C Average 40.75 40.95 40.77 24.12 34.16 46.25
Table 16: mIoU comparison of MLMP and baseline methods on the P59 dataset, evaluated on both the original images and 15 corruption types.
OVSS Backbone: NACLIP Adaptation Method
Dataset: P59 No Adapt. TENT TPT WATT CLIPArTT MLMP
Original 28.23 28.73 $\pm$0.02 28.26 $\pm$0.01 16.55 $\pm$0.05 24.60 $\pm$0.00 31.95 $\pm$0.02
Gaussian noise 21.53 21.49 $\pm$0.01 21.59 $\pm$0.01 11.27 $\pm$0.04 16.42 $\pm$0.03 24.84 $\pm$0.00
Shot noise 22.35 22.31 $\pm$0.01 22.26 $\pm$0.00 11.80 $\pm$0.04 17.47 $\pm$0.00 25.62 $\pm$0.01
Impulse noise 21.74 21.59 $\pm$0.01 21.72 $\pm$0.00 11.24 $\pm$0.03 17.17 $\pm$0.00 24.75 $\pm$0.01
Defocus blur 25.42 25.14 $\pm$0.00 25.41 $\pm$0.00 14.31 $\pm$0.04 20.71 $\pm$0.00 27.95 $\pm$0.03
Glass blur 25.03 24.70 $\pm$0.01 25.01 $\pm$0.00 13.93 $\pm$0.03 20.63 $\pm$0.00 27.66 $\pm$0.05
Motion blur 26.11 26.00 $\pm$0.02 26.13 $\pm$0.01 15.13 $\pm$0.03 21.72 $\pm$0.00 29.12 $\pm$0.03
Zoom blur 19.20 19.49 $\pm$0.03 19.20 $\pm$0.00 10.86 $\pm$0.03 15.39 $\pm$0.00 21.98 $\pm$0.02
Snow 22.45 22.29 $\pm$0.02 22.45 $\pm$0.00 13.60 $\pm$0.04 19.17 $\pm$0.01 25.47 $\pm$0.01
Frost 21.95 21.94 $\pm$0.00 21.97 $\pm$0.00 12.83 $\pm$0.03 18.77 $\pm$0.04 24.52 $\pm$0.01
Fog 24.85 24.91 $\pm$0.03 24.86 $\pm$0.00 13.85 $\pm$0.05 20.66 $\pm$0.00 27.84 $\pm$0.01
Brightness 27.39 27.80 $\pm$0.01 27.42 $\pm$0.00 15.55 $\pm$0.04 23.56 $\pm$0.00 30.63 $\pm$0.01
Contrast 23.55 23.58 $\pm$0.01 23.54 $\pm$0.00 13.38 $\pm$0.05 19.12 $\pm$0.00 26.95 $\pm$0.00
Elastic transform 23.30 23.86 $\pm$0.02 23.30 $\pm$0.01 12.88 $\pm$0.04 20.28 $\pm$0.00 27.16 $\pm$0.01
Pixelate 27.61 27.55 $\pm$0.01 27.62 $\pm$0.00 15.84 $\pm$0.05 23.47 $\pm$0.00 31.23 $\pm$0.01
JPEG compression 25.75 25.51 $\pm$0.05 25.75 $\pm$0.00 14.09 $\pm$0.03 21.20 $\pm$0.00 29.66 $\pm$0.02
P59-C Average 23.88 23.88 23.88 13.37 19.72 27.03
Table 17: mIoU comparison of MLMP and baseline methods on the P60 dataset, evaluated on both the original images and 15 corruption types.
OVSS Backbone: NACLIP Adaptation Method
Dataset: P60 No Adapt. TENT TPT WATT CLIPArTT MLMP
Original 24.95 25.29 24.98 $\pm$0.01 14.77 $\pm$0.04 21.88 $\pm$0.01 27.99 $\pm$0.03
Gaussian noise 19.31 19.17 $\pm$0.01 19.34 $\pm$0.01 10.29 $\pm$0.03 14.91 $\pm$0.01 $2$$\pm$0.02
Shot noise 19.98 19.83 $\pm$0.02 19.91 $\pm$0.01 10.82 $\pm$0.03 15.82 $\pm$0.01 22.81 $\pm$0.01
Impulse noise 19.56 19.27 $\pm$0.01 19.54 $\pm$0.00 10.34 $\pm$0.03 15.58 $\pm$0.00 22.26 $\pm$0.01
Defocus blur 22.74 22.38 $\pm$0.01 22.72 $\pm$0.01 12.79 $\pm$0.04 18.67 $\pm$0.00 24.92 $\pm$0.01
Glass blur 22.49 22.05 $\pm$0.01 22.48 $\pm$0.01 12.64 $\pm$0.03 18.61 $\pm$0.00 24.71 $\pm$0.05
Motion blur 23.28 23.01 $\pm$0.03 23.30 $\pm$0.01 13.53 $\pm$0.03 19.51 $\pm$0.01 25.71 $\pm$0.03
Zoom blur 17.42 17.58 $\pm$0.03 17.42 $\pm$0.00 9.95 $\pm$0.03 14.07 $\pm$0.01 19.70 $\pm$0.02
Snow 20.06 19.80 $\pm$0.01 20.06 $\pm$0.01 12.29 $\pm$0.04 17.20 $\pm$0.02 22.64 $\pm$0.02
Frost 19.71 19.61 $\pm$0.01 19.73 $\pm$0.01 11.59 $\pm$0.03 17.02 $\pm$0.04 22.02 $\pm$0.01
Fog 22.20 22.09 $\pm$0.02 22.20 $\pm$0.01 12.50 $\pm$0.03 18.53 $\pm$0.01 24.89 $\pm$0.00
Brightness 24.35 24.56 $\pm$0.01 24.37 $\pm$0.01 13.87 $\pm$0.04 21.05 $\pm$0.01 27.02 $\pm$0.00
Contrast 21.09 21.00 $\pm$0.03 21.08 $\pm$0.01 12.04 $\pm$0.03 17.24 $\pm$0.01 24.17 $\pm$0.02
Elastic transform 21.17 21.46 $\pm$0.01 21.19 $\pm$0.01 11.70 $\pm$0.04 18.53 $\pm$0.01 24.27 $\pm$0.02
Pixelate 24.52 24.33 $\pm$0.01 24.54 $\pm$0.01 14.17 $\pm$0.04 21.01 $\pm$0.01 27.48 $\pm$0.00
JPEG compression 22.99 22.65 $\pm$0.02 24.54 $\pm$0.01 12.65 $\pm$0.03 19.05 $\pm$0.01 26.24 $\pm$0.01
P60-C Average 21.39 21.25 21.49 12.08 17.79 24.07
Table 18: mIoU comparison of MLMP and baseline methods on the CityScapes dataset, evaluated on both the original images and 15 corruption types.
OVSS Backbone: NACLIP Adaptation Method
Dataset: CityScapes No Adapt. TENT TPT WATT MLMP
Original 29.49 30.54 $\pm$0.04 29.57 $\pm$0.01 20.77 $\pm$0.02 33.35 $\pm$0.03
Gaussian noise 16.14 15.00 $\pm$0.02 15.94 $\pm$0.01 10.19 $\pm$0.01 15.32 $\pm$0.00
Shot noise 20.18 19.38 $\pm$0.01 20.15 $\pm$0.01 12.41 $\pm$0.02 20.57 $\pm$0.04
Impulse noise 16.37 14.59 $\pm$0.02 16.35 $\pm$0.01 8.56 $\pm$0.02 15.76 $\pm$0.05
Defocus blur 23.34 23.50 $\pm$0.02 23.46 $\pm$0.01 15.63 $\pm$0.02 24.86 $\pm$0.08
Glass blur 24.41 24.04 $\pm$0.04 24.29 $\pm$0.00 15.44 $\pm$0.01 25.70 $\pm$0.02
Motion blur 24.73 24.98 $\pm$0.00 24.91 $\pm$0.01 14.21 $\pm$0.03 26.02 $\pm$0.01
Zoom blur 14.08 14.57 $\pm$0.06 14.08 $\pm$0.01 5.85 $\pm$0.02 14.04 $\pm$0.04
Snow 19.88 20.14 $\pm$0.01 19.90 $\pm$0.01 9.27 $\pm$0.03 22.23 $\pm$0.03
Frost 16.78 16.47 $\pm$0.02 16.78 $\pm$0.01 10.69 $\pm$0.01 17.12 $\pm$0.07
Fog 22.47 22.87 $\pm$0.05 22.25 $\pm$0.01 14.24 $\pm$0.01 23.98 $\pm$0.03
Brightness 28.45 29.30 $\pm$0.01 28.47 $\pm$0.01 19.99 $\pm$0.02 31.88 $\pm$0.04
Contrast 16.10 16.35 $\pm$0.03 16.07 $\pm$0.01 11.05 $\pm$0.01 17.04 $\pm$0.00
Elastic transform 29.14 30.16 $\pm$0.02 29.02 $\pm$0.01 19.99 $\pm$0.03 32.86 $\pm$0.00
Pixelate 28.67 29.59 $\pm$0.01 28.69 $\pm$0.01 18.30 $\pm$0.01 32.72 $\pm$0.04
JPEG compression 23.69 23.62 $\pm$0.01 23.67 $\pm$0.01 15.98 $\pm$0.02 25.26 $\pm$0.03
CityScapes-C Average 21.63 21.64 21.60 13.45 23.02